OpenStudy (anonymous):

change the given angles to equal angles expressed in decimal form to the neareast 0.001 degree. -107 degree 16' 23"

OpenStudy (anonymous):

i got -107.237 is this correct?

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

\[ -\left(107+\frac{16}{60}+\frac{23}{60^2}\right) \]

OpenStudy (anonymous):

okay got it